# RayGauss: Volumetric Gaussian-Based Ray Casting for Photorealistic Novel View Synthesis

### v1

@@ Abstract
Our work closes this gap by providing a physically consistent formulation of the emitted radiance $c$ and density $\sigma$, decomposed with Gaussian functions associated with Spherical Gaussians/Harmonics for all-frequency colorimetric representation. We also introduce a method enabling differentiable ray casting of irregularly distributed Gaussians using an algorithm that integrates radiance fields slab by slab and leverages a BVH structure. This allows our approach to finely adapt to the scene while avoiding splatting artifacts. As a result, we achieve superior rendering quality compared to the state-of-the-art while maintaining reasonable training times and achieving inference speeds of 25 FPS on the Blender dataset.
Project page with videos and code: raygauss.github.io
@@ 1 Introduction
Novel View Synthesis (NVS), which combines existing views of a scene to generate images from unknown viewpoints, saw significant advancements with the publication of Neural Radiance Fields (NeRF) [20]. This novel approach leverages a differentiable physical rendering algorithm [18], to learn radiance fields through a neural network, enabling the generation of high-quality photorealistic images. Since then, several works aimed at improving NeRF’s performance, some focusing on finding more efficient representations of radiance fields in terms of computation time and rendering quality. Various data structures were investigated to substitute the neural network and represent the radiance fields: voxel grids [7][29][21], sets of tetrahedra [16], point clouds[31][11]. These structures have the advantage of enabling the representation of radiance fields using local parameters, thereby reducing computation times by limiting field evaluation to local parameters instead of relying on a neural network representing the entire scene. These methods also aim to enhance rendering quality by allowing a more precise representation of radiance fields through data structures better suited to the scene’s geometry.
@@
Our work follows this approach by proposing a novel definition of radiance fields based on their decomposition into a weighted sum of locally supported elliptical basis functions with optimizable positions, orientations, and scales. Unlike voxel grid-based methods, this allows for finer adaptation to the scene’s geometry without limiting representation resolution. Furthermore, a combination of spherical harmonics (SH) and spherical Gaussians (SG) is associated with these basis functions to better represent high-frequency color variations such as specular reflections. Additionally, we leverage the volume ray casting approach, which generates high-quality images with less restrictive simplifications compared to splatting-based approaches like 3D Gaussian Splatting [11]. Indeed, splatting algorithms simplify the scene using multiple approximations, allowing faster rendering but causing many artifacts when deviating from the training views, such as flickering due to a sudden change in primitives ordering[25]. For these reasons, our approach achieves higher rendering quality by combining the best of both worlds: Firstly, radiance fields are represented by elliptical functions associated with efficient SH/SG radiance parameters that adapt very well to the scene geometry and appearance. Secondly, the Volume Raycasting algorithm avoids the artifacts of Splatting, such as flickering.
@@
Furthermore, implementing the Raycasting algorithm with sparse primitives is challenging because evaluating the radiance fields at a given point requires knowing which primitives/basis functions contribute to the calculations, which is much less straightforward than when using a voxel grid, for example. Therefore, we introduce a specialized ray casting algorithm for sparse basis functions that sequentially integrates the radiance field slab-by-slab, meaning it accumulates colorimetric properties along the ray in slab of space corresponding to multiple samples, rather than processing individual samples[13]. The implementation of this algorithm and its backpropagation relies on a Bounding Volume Hierarchy (BVH) implemented on GPU with the OptiX library. This allows interactive rendering times and reasonable training times. Thus, our contributions are as follows:
@@
  • Volume Ray Casting of radiance fields decomposed into elliptical basis functions combined with SH/SG for efficient adaptation to scene geometry and appearance.
@@ 2 Background and Related Work
Thus, NeRF’s contribution is using a differentiable volumetric rendering algorithm to learn the scene parameters represented by a neural network. Subsequent research papers have proposed improvements to the scene representation by using different structures to store scene parameters locally, thus eliminating the dependency on a computationally costly neural network representing the entire scene and improving the appearance by better adapting the local parameters than a global MLP. Moreover, due to their sparse nature, some data structures can apply classic ray-casting acceleration strategies such as empty space skipping [8]. Among these structures, one can notably mention voxel grids [29], grids with hash encoding [21], grids of small MLPs [26], sets of tetrahedrons [16], and point clouds [31]. Our hypothesis is that point clouds are best suited to adapt to scenes. Each point can host a locally supported function, such as Gaussian functions [11], representing local scene details. Optimizing these functions’ positions and shapes enables precise adaptation to scene geometry, unlike voxel grids constrained by their resolution. Furthermore, to the best of our knowledge, no method currently allows for optimizing a scene represented only by irregularly spaced basis functions with the ray-casting algorithm without coarse approximations. The approach in Point-NeRF [31] is interesting but requires using a MVSNet, which proves cumbersome. Additionally, its expression of $c$ and $\sigma$ involves using a neural network and only considers the eight closest neighbors in its calculations. Moreover, as discussed in the following section, the 3D Gaussian Splatting method and numerous point cloud rendering methods [33] make the splatting approximation, which can cause visible artifacts such as flickering when changing viewpoints. This choice is due to the scarcity of algorithms that efficiently perform ray casting on local basis functions associated with points.
@@
Where $x^{\prime}$ the pixel coordinates in the camera frame, $\mu^{\prime}_{i}$ and $\Sigma^{\prime}_{i}$ are the mean and covariance of the Gaussian projected into the camera frame, and $\mathcal{N}$ denotes the set of Gaussians. Interested readers can find the implementation details in the original paper [11] and the following review [5]. Like the discretized equation of ray casting, the formulation of splatting 5 is also an approximation of Eq. 3. However, the approximations used differ: Gaussians are assumed to be non-overlapping, which does not hold for complex surfaces. Indeed, multiple overlapping Gaussians are required for precise scene geometry reconstruction. Then, each Gaussian contributes at most once per ray, aggregated at a distance corresponding to the depth of the Gaussian (Gaussian mean projected along the viewing direction). Also, the sequential addition of Gaussian contributions based on global depth is an approximation. It does not account for the proper intersection between the Gaussians and a given ray, further reducing the coherence of the representation. Finally, the projection of Gaussians into the camera frame is also an approximation introduced in [34]. Thus, in our approach, we use the ray casting algorithm to suffer fewer approximations because we hypothesize that, given a high enough sampling rate, it maintains better coherence in scene representation during training and inference and should yield better graphical results.
@@ 3 Scene Representation
$\sigma(\mathbf{x})$ $=\sum_{l=1}^{N}w_{\sigma_{l}}\phi_{l}(\mathbf{x})$ (6)
$c(\mathbf{x},\mathbf{d})$ $=\sum_{l=1}^{N}w_{c_{l}}(\mathbf{d})\psi_{l}(\mathbf{x})$
@@ 3.1 Irregularly Spatially-Distributed Emissive and Absorptive Primitives
$c(\mathbf{x},\mathbf{d})=\dfrac{\sum_{l=1}^{N}c_{l}({\mathbf{d}})\tilde{\sigma_{l}}\phi_{l}(\mathbf{x})}{\sum_{l=1}^{N}\tilde{\sigma_{l}}\phi_{l}(\mathbf{x})}$ (11)
@@ 3.2 Selection of the basis function
Where $\mathbf{\mu}$ is the mean vector, $\mathbf{\Sigma}$ is the covariance matrix. Here, we can notice that $\sqrt{(\mathbf{x}-\mathbf{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{\mu})}$ is the Mahalanobis distance associated with the matrix $\Sigma$. To achieve a more suitable parameterization, the covariance matrix $\mathbf{\Sigma}$, being positive and symmetric, can be decomposed according to the spectral theorem into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$ such that:
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$
Additionally, the rotation matrix can be expressed in terms of quaternions $\mathbf{q}\in\mathbb{R}^{4}$ such that $\mathbf{R}=\mathbf{R}(\mathbf{q})$, and the scaling matrix $\mathbf{S}$ can be expressed in terms of $\mathbf{s}=(s_{x},s_{y},s_{z})$ such that $\mathbf{S}=\text{diag}(s_{x},s_{y},s_{z})$. Thus, we will refer to our Gaussian basis functions as $\mathcal{G}(\mathbf{x};\mathbf{\mu},\mathbf{q},\mathbf{s})$.
Therefore, we can now describe our density field $\sigma(\mathbf{x})$ and emitted radiance field $c(\mathbf{x},\mathbf{d})$ as follows:
@@
$c(\mathbf{x},\mathbf{d})$ $=\frac{\sum_{l=1}^{N}c_{l}({\mathbf{d}})\tilde{\sigma_{l}}\mathcal{G}(\mathbf{x};\mu_{l},\mathbf{q}_{l},\mathbf{s}_{l})}{\sum_{l=1}^{N}\tilde{\sigma_{l}}\mathcal{G}(\mathbf{x};\mu_{l},\mathbf{q}_{l},\mathbf{s}_{l})}$
@@ 3.3 Direction-dependent emitted radiance
To correctly represent objects whose color appearance varies according to the direction of observation, we made the emitted radiance of our primitives dependent on the viewing direction. One parameterization that gave the best results was a combination of spherical harmonics and spherical Gaussians, such that emitted radiance of l-th primitive is:
@@
This choice is motivated by the fact that low-degree Spherical Harmonics are better suited to represent low-frequency variations. In contrast, Spherical Gaussians can represent any frequency, including high frequencies. Thus, their combined use allows for representing both low and high-frequency phenomena. Similar considerations can be found for different applications in articles such as MixLight [10].
@@ 4.1 Scene initialization
To initialize this representation, we leverage sparse point clouds obtained via Structure from Motion (SfM), where the position and color of these points are used to initialize colorimetric parameters and the center of the basis functions $\mu_{l}$. Additionally, the quaternions are initialized to match the identity matrix, and the scale parameters are initialized to the average distance from each point to its three nearest neighbors.
@@ 4.2 Scene Optimization
Once the representation is initialized, we optimize it using a stochastic gradient descent algorithm. To do this, we rely on a set of training images, where each image is associated with the pose of its camera. Iteratively, over the training set, we generate the scene image from the current camera pose and compare it to the current training image using a given loss function. Subsequently, we compute gradients of the loss function with respect to the scene parameters, allowing us to update them by backpropagation. In particular, the loss function $L$ is a weighted sum of the $L_{1}$ loss function and the structural dissimilarity (DSSIM):
@@
where $\lambda$ is a weighting factor between the L1 loss and DSSIM, we follow best practices developed in 3D Gaussian Splatting [11]. Furthermore, we are in a similar case to 3D Gaussian Splatting regarding adaptive control of primitives. Indeed, we also deal with primitives irregularly distributed in space. For this reason, we rely on the point cloud densification heuristics described in [11]. Regarding removing outlier points, we eliminate points whose density falls below a fixed threshold $\sigma_{\epsilon}$. Additionally, we progressively unlock the ability to learn different colorimetric parameters to enforce model coherence at low frequencies before moving to higher frequencies. Thus, we gradually unlock the degrees of spherical harmonics and then unlock spherical Gaussians.
@@ 5 Spatially irregularly distributed base functions ray casting
This approach has been scarcely explored due to the major obstacle of implementing an algorithm that enables efficient differentiable volume ray casting for a scene defined with irregularly distributed basis functions. Firstly, it should be noted that since the support of our basis functions is unbounded in the chosen case (Gaussian functions), it would potentially require evaluating all basis functions to compute $\sigma$ and $c$ at any spatial position. This promises to slow down the rendering of our images significantly. To mitigate this, we truncate our basis functions to a local domain. The criterion retained to determine this domain for each primitive is as follows:
@@ 6.2 Results and Evaluation
To quantify our method’s performance, we tested it on two widely used datasets in the NVS domain: the Blender dataset [20] and the Mip-NeRF 360°dataset [2]. This allows us to test our method across a wide range of situations. The Blender dataset contains 8 scenes with complex objects exhibiting non-Lambertian effects. Each scene consists of 100 training images, 200 test images, and the corresponding exact camera parameters for 360°$\degree$ views of the object. The Mip NeRF 360°$\degree$ dataset comprises 9 scenes, 5 outdoor and 4 indoor, each featuring a central object surrounded by a complex background. Each scene is accompanied by images with camera parameters estimated by COLMAP [27]. We evaluate our method by downscaling images by a factor of 8 for this dataset due to large image size. Moreover, we compare our results primarily with 3D Gaussian Splatting [11] and subsequent research papers such as Mip-Splatting [32] due to the similarity in the representation used.
Results on the Blender dataset: For synthetic scenes in the Blender dataset, we provide quantitative comparisons in Tab. 1 and qualitative comparisons in Fig. 4 with state-of-the-art methods. For the results in Tab. 1, the values used are those reported in the state-of-the-art methods, except for those denoted with an *, which have been recalculated using the available code.
Quantitative results show that our approach surpasses state-of-the-art results on most individual scenes. Additionally, we achieve an average PSNR of 34.34 dB, attesting our model’s ability to finely adapt to the geometry and appearance of the scenes compared to other state-of-the-art methods. Qualitative results also support this, for example, with a boat hull better reconstructed in the ship scene compared to 3D Gaussian Splatting reconstruction (Figure 4).
@@
Results on the Mip-NeRF 360 Dataset: Regarding Mip-NeRF 360, we provide quantitative results in Tab. 2 and qualitative comparisons in Fig. 2. A general trend we observe is that our approach performs better on most indoor scenes than current state-of-the-art methods such as Mip-Splatting [2] and slightly worse on outdoor scenes. This may be due to the adaptive control of gaussians strategy, which adapts well in the case of Gaussian Splatting[11] but may perform worse with our approach. Nonetheless, we achieved better overall results across all scenes with an average PSNR of 29.96 dB, which is +0.81 dB higher than Gaussian Splatting.
@@
Performance Analysis: We discuss here results in Tab. 3, which show the performance of our approach compared to Gaussian Splatting. These metrics were obtained using an RTX 4090 GPU with 24GB of RAM on 800x800 pixel images. Our approach has slower training and inference times compared to Gaussian Splatting due to the computational cost of ray casting. Generally, our computation times depend predominantly on the number of rays, while Gaussian Splatting’s times are influenced by the number of projected primitives. However, our training times remain reasonable (30 minutes per scene on the Blender dataset) and real-time inference is achieved on Blender, although slower on Mip-NeRF 360. We also note that our method generates fewer primitives for better visual quality. Thus, Gaussian Splatting likely adds more Gaussians to compensate for splatting artifacts.

### v2

@@ Abstract
Our work closes this gap by providing a physically consistent formulation of the emitted radiance $c$ and density $\sigma$, decomposed with Gaussian functions associated with Spherical Gaussians/Harmonics for all-frequency colorimetric representation. We also introduce a method enabling differentiable ray casting of irregularly distributed Gaussians using an algorithm that integrates radiance fields slab by slab and leverages a BVH structure. This allows our approach to finely adapt to the scene while avoiding splatting artifacts. As a result, we achieve superior rendering quality compared to the state-of-the-art while maintaining reasonable training times and achieving inference speeds of 25 FPS on the Blender dataset. The associated code is available at: github.com/hugobl1/ray_gauss.
@@ 1 Introduction
Novel View Synthesis (NVS), which combines existing views of a scene to generate images from unknown viewpoints, saw significant advancements with the publication of Neural Radiance Fields (NeRF) [21]. This novel approach leverages a differentiable physical rendering algorithm [19], to learn radiance fields through a neural network, enabling the generation of high-quality photorealistic images. Since then, several works aimed at improving NeRF’s performance, some focusing on finding more efficient representations of radiance fields in terms of computation time and rendering quality. Various data structures were investigated to substitute the neural network and represent the radiance fields: voxel grids [8][30][22], sets of tetrahedra [17], point clouds[32][12], hybrid representations[6]. These structures have the advantage of enabling the representation of radiance fields using local parameters, thereby reducing computation times by limiting field evaluation to local parameters instead of relying on a neural network representing the entire scene. These methods also aim to enhance rendering quality by allowing a more precise representation of radiance fields through data structures better suited to the scene’s geometry.
@@
Our work follows this approach by proposing a novel definition of radiance fields based on their decomposition into a weighted sum of locally supported elliptical basis functions with optimizable positions, orientations, and scales. Unlike voxel grid-based methods, this allows for finer adaptation to the scene’s geometry without limiting representation resolution. Furthermore, a combination of spherical harmonics (SH) and spherical Gaussians (SG) is associated with these basis functions to better represent high-frequency color variations such as specular reflections. Additionally, we use volume ray casting, which produces high-quality images with fewer simplifications than splatting-based methods like 3D Gaussian Splatting[12]. Indeed, splatting algorithms simplify the scene using multiple approximations, allowing faster rendering but causing many artifacts when deviating from the training views, such as flickering due to a sudden change in primitives ordering[26]. For these reasons, our approach achieves higher rendering quality by combining the best of both worlds: Firstly, radiance fields are represented by elliptical functions associated with efficient SH/SG radiance parameters that adapt very well to the scene geometry and appearance. Secondly, the Volume Raycasting algorithm avoids the artifacts of Splatting, such as flickering.
@@
Furthermore, implementing the Raycasting algorithm with sparse primitives is challenging because evaluating the radiance fields at a given point requires knowing which primitives/basis functions contribute to the calculations, which is much less straightforward than when using a voxel grid, for example. Therefore, we introduce a ray casting algorithm for sparse basis functions that sequentially integrates the radiance field slab-by-slab, meaning it accumulates colorimetric properties along the ray in slab of space corresponding to multiple samples, rather than processing individual samples[14]. The implementation of this algorithm and its backpropagation relies on a Bounding Volume Hierarchy (BVH) implemented on GPU with the OptiX library. This allows interactive rendering times and reasonable training times. Thus, our contributions are as follows:
@@
  • Volume Ray Casting of radiance fields decomposed into elliptical basis functions combined with SH/SG for efficient adaptation to geometry and appearance.
@@ 2 Background and Related Work
NeRF’s key contribution is using a differentiable volumetric rendering algorithm to learn scene parameters with a neural network. Later research improved scene representation by using different structures to locally store parameters, reducing reliance on a costly global neural network and enhancing appearance through better local parameter adaptation. Moreover, due to their sparse nature, some data structures can apply classic ray-casting acceleration strategies such as empty space skipping [9]. Among these structures, one can notably mention voxel grids [30], grids with hash encoding [22], grids of small MLPs [27], sets of tetrahedrons [17], and point clouds [32]. Our hypothesis is that point clouds are best suited to adapt to scenes. Each point can host a locally supported function, such as Gaussian functions [12], representing local scene details. Optimizing these functions’ positions and shapes enables precise adaptation to scene geometry, unlike voxel grids constrained by their resolution. Furthermore, to the best of our knowledge, no method currently allows for optimizing a scene represented only by irregularly spaced basis functions with the ray-casting algorithm without coarse approximations. The approach in Point-NeRF [32] is interesting but requires using a MVSNet, which proves cumbersome. Additionally, its expression of $c$ and $\sigma$ involves using a neural network and only considers the eight closest neighbors in its calculations. The NeuRBF approach [6] is also promising, as it uses a hybrid representation with adaptive RBFs for fine details and grid-based RBFs for efficient continuous representation. The use of grid-based RBFs is needed because the algorithm relies on a limited number of nearest adaptive RBFs (5 for Blender dataset), causing discontinuities. Like Point-NeRF, this limits its ability to fully exploit the approximation power of sparse basis functions. Moreover, as discussed in the following section, 3D Gaussian Splatting and numerous point cloud rendering methods [34] make the splatting approximation, which can cause visible artifacts such as flickering when changing viewpoints. This choice is due to the scarcity of algorithms that efficiently perform ray casting on local basis functions associated with points.
@@
Where $x^{\prime}$ the pixel coordinates in the camera frame, $\mu^{\prime}_{i}$ and $\Sigma^{\prime}_{i}$ are the mean and covariance of the Gaussian projected into the camera frame, and $\mathcal{N}$ denotes the set of Gaussians. More implementation details are available in the original paper [12] and the following review [5]. Like the discretized equation of ray casting, the splatting formulation 5 is also an approximation of Eq. 3. However, the approximations used differ: Gaussians are assumed to be non-overlapping, which does not hold for complex surfaces. Indeed, multiple overlapping Gaussians are required for precise scene geometry reconstruction. Then, each Gaussian contributes at most once per ray, aggregated at a distance corresponding to the depth of the Gaussian (Gaussian mean projected along the viewing direction). Also, the sequential addition of Gaussian contributions based on global depth is an approximation. It does not account for the proper intersection between the Gaussians and a given ray, further reducing the coherence of the representation. Finally, the projection of Gaussians into the camera frame is also an approximation introduced in [35]. Thus, we use the ray casting algorithm to suffer fewer approximations because we hypothesize that, given a high enough sampling rate, it maintains better coherence in scene representation during training and inference and should yield better graphical results.
@@ 3 Scene Representation
$\sigma(\mathbf{x})$ $=\sum_{l=1}^{N}\tilde{\sigma_{l}}\phi_{l}(\mathbf{x})$ (6)
$c(\mathbf{x},\mathbf{d})$ $=\sum_{l=1}^{N}c_{l}({\mathbf{d}})\psi_{l}(\mathbf{x})$
@@ 3.1 Irregularly Spatially-Distributed Emissive and Absorptive Primitives
$c(\mathbf{x},\mathbf{d})=\dfrac{\sum_{l=1}^{N}c_{l}({\mathbf{d}})\tilde{\sigma_{l}}\phi_{l}(\mathbf{x})}{\sum_{m=1}^{N}\tilde{\sigma}_{m}\phi_{m}(\mathbf{x})}$ (11)
@@ 3.2 Selection of the basis function
Where $\mathbf{\mu}$ is the mean vector, $\mathbf{\Sigma}$ is the covariance matrix. Here, we can notice that $\sqrt{(\mathbf{x}-\mathbf{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{\mu})}$ is the Mahalanobis distance associated with the matrix $\Sigma$. To achieve a more suitable parameterization, the covariance matrix $\mathbf{\Sigma}$, being positive and symmetric, will be decomposed according to the spectral theorem into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$ such that $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$. The rotation $\mathbf{R}$, as part of the covariance matrix $\mathbf{\Sigma}$ decomposition, can be expressed in terms of unit quaternions $\mathbf{q}\in\mathbb{R}^{4}$ such that $\mathbf{R}=\mathbf{R}(\mathbf{q})$, and the scaling matrix $\mathbf{S}$ can be expressed in terms of $\mathbf{s}=(s_{x},s_{y},s_{z})$ such that $\mathbf{S}=\text{diag}(s_{x},s_{y},s_{z})$. Thus, we will refer to our Gaussian basis functions as $\mathcal{G}(\mathbf{x};\mathbf{\mu},\mathbf{q},\mathbf{s})$. We can now describe our density field $\sigma(\mathbf{x})$ and emitted radiance field $c(\mathbf{x},\mathbf{d})$ as follows:
@@
$c(\mathbf{x},\mathbf{d})$ $=\frac{\sum_{l=1}^{N}c_{l}({\mathbf{d}})\tilde{\sigma}_{l}\mathcal{G}(\mathbf{x};\mu_{l},\mathbf{q}_{l},\mathbf{s}_{l})}{\sum_{m=1}^{N}\tilde{\sigma}_{m}\mathcal{G}(\mathbf{x};\mu_{m},\mathbf{q}_{m},\mathbf{s}_{m})}$
@@ 3.3 Direction-dependent emitted radiance
To correctly represent objects whose color appearance varies by observation direction, we made the emitted radiance of our primitives dependent on the viewing direction. One parameterization that gave the best results was a combination of spherical harmonics and spherical Gaussians, such that emitted radiance of l-th primitive is:
@@
This choice is motivated by low-degree Spherical Harmonics being better suited to represent low-frequency variations. In contrast, Spherical Gaussians can represent any frequency, including high frequencies. Thus, their combined use allows for representing both low and high-frequency phenomena. Similar considerations can be found for different applications in articles such as MixLight [11].
@@ 4.1 Scene initialization
To initialize this representation, we leverage sparse point clouds obtained via Structure from Motion (SfM), where the position and color of these points are used to initialize colorimetric parameters and the center of the basis functions $\mu_{l}$. Additionally, unit quaternions are initialized to match the identity matrix, and the scale parameters are initialized to the average distance from each point to its three nearest neighbors.
@@ 4.2 Scene Optimization
After initialization, we optimize the representation using stochastic gradient descent, relying on training images paired with their camera poses. Iteratively, over the training set, we generate the scene image from the current camera pose and compare it to the current training image using a given loss function. We then compute loss gradients with respect to scene parameters for backpropagation updates. In particular, the loss function $L$ is a weighted sum of the $L_{1}$ loss function and the structural dissimilarity (DSSIM):
@@
where $\lambda$ is a weighting factor between the L1 loss and DSSIM, we follow best practices developed in 3D Gaussian Splatting [12]. Furthermore, we are in a similar case to 3D Gaussian Splatting regarding adaptive control of primitives. Indeed, we also deal with primitives irregularly distributed in space. For this reason, we rely on the point cloud densification heuristics described in [12]. Additionally, we remove outlier points with density below a fixed threshold $\sigma_{\epsilon}$ and progressively unlock colorimetric parameters, starting with spherical harmonics from low to high degree and then spherical Gaussians, to enforce model coherence from low to high frequencies.
@@ 5 Spatially irregularly distributed base functions ray casting
This approach has been scarcely explored due to the major obstacle of implementing an algorithm that enables efficient differentiable volume ray casting for a scene defined with irregularly distributed basis functions. Firstly, since the support of our basis functions is unbounded in the chosen case (Gaussian functions), it would potentially require evaluating all basis functions to compute $\sigma$ and $c$ at any spatial position. This promises to slow down the rendering of our images significantly. To mitigate this, we truncate our basis functions to a local domain. The criterion retained to determine this domain for each primitive is as follows:
@@ 6.2 Results and Evaluation
To quantify our method’s performance, we tested it on two widely used NVS datasets: the Blender dataset [21] and the Mip-NeRF 360°dataset [2]. This allows us to test our method across a wide range of situations. The Blender dataset includes 8 scenes with complex objects exhibiting non-Lambertian effects, each with 100 training images, 200 test images, and exact camera parameters for 360°$\degree$ views. The Mip NeRF 360°$\degree$ dataset comprises 9 scenes: 5 outdoor and 4 indoor, each featuring a central object with a complex background. Each scene includes images with camera parameters estimated by COLMAP [28]. We evaluate our method by downscaling images by a factor of 8 for this dataset due to large image size. Moreover, we compare our results mainly with 3D Gaussian Splatting [12] and subsequent research papers such as Mip-Splatting [33] due to the similarity in the representation. The results in Tables 1 and 2 are from state-of-the-art methods, except those marked with *, which were recalculated using available code.
Results on the Blender dataset: For synthetic scenes in the Blender dataset, we provide quantitative comparisons in Tab. 1 and qualitative comparisons in Fig. 4 with state-of-the-art methods. Quantitative results show that our approach surpasses state-of-the-art results on most individual scenes. Additionally, our average PSNR of 34.53 dB highlights our model’s superior adaptation to scene geometry and appearance compared to other state-of-the-art methods. Qualitative results also support this, for example, with a boat hull better reconstructed in the ship scene compared to 3D Gaussian Splatting reconstruction (Figure 4). NeuRBF [6] performs well, similar to our method, but has an inference time of a few seconds per image—about 100x longer than ours (25 FPS on Blender dataset, see Tab. 3)—and does not support the Mip-NeRF360 dataset.
@@
Results on the Mip-NeRF 360 Dataset: Regarding Mip-NeRF 360, we provide quantitative results in Tab. 2 and qualitative comparisons in Fig. 2. A general trend is that our approach performs better on most indoor scenes than state-of-the-art methods like Mip-Splatting [2], and slightly worse on outdoor scenes. This may be due to the adaptive control of gaussians strategy, which adapts well in Gaussian Splatting [12] but may perform worse with our approach. Nevertheless, we achieved better overall results than Mip-Splatting and 3D Gaussian Splatting, with an average PSNR of 29.71 dB, +0.56 dB higher than 3D Gaussian Splatting. Zip-NeRF performs slightly better on Mip-NeRF 360 but requires over 7 hours of training per scene and about 1.5 seconds per image for inference, which is significantly higher than our approach (see Tab. 3).
@@
Performance Analysis: We discuss here results in Tab. 3, which show the performance of our approach compared to Gaussian Splatting. These metrics were obtained using an RTX 4090 GPU with 24GB of RAM on 800x800 pixel images. Our approach has slower training and inference times compared to Gaussian Splatting due to the computational cost of ray casting. Generally, our computation times depend mainly on the number of rays, while Gaussian Splatting’s times are influenced by the number of projected primitives. However, our training times remain reasonable (30 minutes per scene on Blender) and real-time inference is achieved on Blender, although slower on Mip-NeRF 360. We also note that our method generates fewer primitives for better visual quality. Thus, Gaussian Splatting likely adds more Gaussians to compensate for splatting artifacts.